OpenStudy (anonymous):

Trevor's dad decided to get him a cell phone for his birthday. Cell Plus has a plan that costs $25.00 per month plus an additional $0.06 per minute. Cell Best has a plan that costs $30.00 per month plus an additional $0.05 per minute. How many minutes can Trevor talk and both cell plans cost the same amount? Answer 500 100 50 1000 Trevor's dad decided to get him a cell phone for his birthday. Cell Plus has a plan that costs $25.00 per month plus an additional $0.06 per minute. Cell Best has a plan that costs $30.00 per month plus an additional $0.05 per minute. How many minutes can Trevor talk and both cell plans cost the same amount? Answer 500 100 50 1000 @Mathematics

OpenStudy (anonymous):

Has Trevor never heard of unlimited minutes?

OpenStudy (anonymous):

Anyway, here's how you solve it...

OpenStudy (anonymous):

Equation corresponding to Cell Plus Plan: $25.00+($0.06)m=Total Cost Equation corresponding to Cell Best Plan: $30.00+($0.05)m=Total Cost where m is number of minutes used. To find out where they intersect, you set the equations equal to each other and solve for m: $25.00+($0.06)m=$30.00+($0.05)m (0.06-0.05)m=30-25 0.01m=5 m=5/0.01=500 minutes. Hope that helps.